OpenStudy (anonymous):

If \(\color{blue}{g = 10^{100}}\) and \(\color{green}{H = 10^{g}}\) then in which \(\text{Interval}\) does \(\color{red}{g! = 1 \cdot 2 \cdot 3 .......... 10^{100}}\) lie? a) \(10H < g! < H \) b) \(H < g! <10^{H}\) c) \(10H < g! < 10^{H}\) d) \(10^{H} < g!\)

OpenStudy (anonymous):

x=?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Wait, a minute I have to edit question more..

OpenStudy (anonymous):

Isnt x given?

OpenStudy (anonymous):

Not at all..

OpenStudy (anonymous):

If x=1 then option D is correct

OpenStudy (anonymous):

x=1000000000000000000000000000000000000000000000000000000000000000 then none of the option is correct

OpenStudy (anonymous):

What is logic behind it @sauravshakya

OpenStudy (anonymous):

Sorry, I misread, that is not x that is g there..

OpenStudy (unklerhaukus):

option a) isn't right ,

OpenStudy (anonymous):

Hello Uncle Rocks, how are you?? Happy New Year to you and to your Family, First of all.. What is logic behind this question can you teach me??

OpenStudy (anonymous):

Sorry, I was helping someone else, that is why I am late as I usually do..

OpenStudy (unklerhaukus):

hey im alright a) isnt right because \[10H\not <H\]

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

And that is H = 10 raised to the power g and not 9..

OpenStudy (anonymous):

\[\large H = 10^{\text{g}}\]

OpenStudy (unklerhaukus):

I think this is an excellent question for using Stirling's Approximation for large factorials \[\boxed{\log n!=n\log n-n+\mathcal O(\log n)}\] take the log_10 for each side of inequality

OpenStudy (anonymous):

Now, what is Stirling's Approximation? I Have solution for this at the back of book, but I don't want to see it..

OpenStudy (unklerhaukus):

Stirling's Approximation is the thing in the box ^^

OpenStudy (anonymous):

What is O like character there??

OpenStudy (unklerhaukus):

that big O indicates the order of the error in the approximation

OpenStudy (anonymous):

You think differently @UnkleRhaukus For your way of answering this question, I must buy A Mentos..

OpenStudy (unklerhaukus):

the error term is smaller (compared to the approximation) the larger n is , in these examples you can neglect the \(\mathcal O(\cdot)\) term completely

OpenStudy (anonymous):

10H<g! 10*10^g < 1*2*3*...*10^100 10^g < 1*2*3*...*(10^100 -1)*10^(100-1) 10*10*10*...*10 < 1*2*3*...*(10^100 -1) *10^(100-1) Which is TRUE as 0nly first 9 terms of RHS are less than that of LHS So, 10H<g!

OpenStudy (anonymous):

So, it cannot be option B

OpenStudy (anonymous):

g!<10^H log(g!) < log10^H log(1*2*3*...*10^100) < H log1 + log2 +log3 +... +log10^100 < 10^g log1+log2+10g3+...+100 <10*10*10*...*10 This above inequality is TRUE as log1+log2+log3+...+100<100+100+100+...+100<10*10*10*...*10 So, g!<10^H

OpenStudy (anonymous):

Thus, 10H<g!<10^H

OpenStudy (anonymous):

NOTE: